solve the linear system 5x-5y=-20 and 15x +5y=10

Found 2 solutions by Alan3354, TimothyLamb:
Answer by Alan3354(69443)

5x-5y=-20
15x +5y=10
Divide both by 5
---
x - y =-4
3x + y =2
----------- add them

Answer by TimothyLamb(4379)

---
5x - 5y = -20
15x + 5y = 10
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= -0.5
y= 3.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php